3. Calculate and categorize the point elasticity for the (inverse) demand curve P = 24 -20 at P=14, P2=12, Pz=10.
Answer #1

Point elasticity = % change in Qd/% change in price

If Ed>1, elastic, less than 1 inelastic and equal to 1 unit elastic

P

Q= 12-0.5P

% Change in Q

%change in P

Ed = % Change in Qd/% Change in P

Categorize

P0

16

=12-0.5*16 = 4

P1

14

=12-0.5*14 = 5

=(5-4)/4 = 0.25

=(14-16)/16 = -0.125

=0.25/-0.125 = -2

Elastic

P2

12

=12-0.5*12 = 6

=(6-5)/5 = 0.2

=(12-14)/14 = -0.143

=0.2/-0.143 = -1.4

Elastic

P3

10

=12-0.5*10 = 7

=(7-6)/6 = 0.167

=(10-12)/12 = -0.167

=0.167/-0.167 = 1

Unit elastic